OpenStudy (anonymous):

HELP PLEEEASE!!!!!!! FOR A MEDAL!!!!!!!!!!!! Solve the following system of equations. -3x-4y=20 -4x-3y=15

OpenStudy (anonymous):

multiply the top equation by 4 and the bottom equation by 3. You will have -12x in both equations. Then add the two equations and the -12x will cancel out, leaving you with -25y = 125, y=-5. Therefore x=0

OpenStudy (anonymous):

ok so u need to make one of the variables in both equations cancel out

hero (hero):

Now what if we decided to simply multiply both sides of the second equation by -1: Then we would have -3x - 4y = 20 4x + 3y = -15 And after combining the equations together we would have x - y = 5 And upon isolating x: x = 5 + y We could then substitute that into the second equation: 4(5 + y) + 3y = -15 20 + 4y + 3y = -15 20 + 7y = -15 7y = -35 y = -5 And upon back substituting we would solve for x: 4x + 3(-5) = -15 4x - 15 = -15 4x = 15 - 15 4x = 0 x = 0